Check whether the given fractions are equivalent:
 5 / 9, 30 / 54

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Given 5 / 9 and 30 / 54
We have 5× 54 = 270
9 × 30 = 270
5 × 54 = 270
5 × 54 = 9 × 3
Hence, 5 / 9 and 30 / 54 are equivalent fractions

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Let's check whether the following fractions are equivalent,
We have ,
5/9 = 30/54
(5×54) = (30×9)
270 = 270
 
Hence, the given fractions are equivalent.
 
Hope it helps!
Thank you ✌️